# Backpropagation-Free Test-Time Adaptation via Probabilistic Gaussian Alignment

### v5

@@ 4.2 Main Results
Online CLIP CLIP 66.74 - 8m 0.79
TPT TPT_NIPS22 68.95 2.21 9h 45m 4.29
DiffTPT DiffTPT_ICCV23 70.30 3.56 > 20h 4.60
TDA TDA_CVPR24 69.51 2.77 50m 0.84
TPS TPS_WACV25 70.38 3.64 1h 19m 1.71
@@
Trans. GDA-CLIP  hard_ICLR24 64.13 -2.61 1.31m 10.03
TransCLIP TransCLIP_NIPS24 70.30 3.56 1.34m 16.17
StatA StatA_CVPR25 69.90 3.16 1.5m 20.74
@@
Vanilla CLIP 64.70 27.52 67.43
Ensemble CLIP 66.56 28.54 67.74
CLIP Template CLIP 66.51 28.44 67.62
GPT  AWT_NIPS24 66.53 28.56 70.76
@@ 4.3 Ablation Studies and Further Analysis
We study how different initializations of the mean $\hat{\boldsymbol{\mu}}$ in Eq. (9) affect performance. We compare 6 types of text-derived prototypes: CLIP’s vanilla template, 7 generic templates, 80 ImageNet-specific templates, GPT-generated descriptions in AWT_NIPS24 , and its variants. As shown in Table 9, diverse and descriptive initializations improve performance, though ADAPT remains robust across all settings. We use GPT-based descriptions in our main results.
@@ A.2 Online Test-time Distribution Estimation
For the third term, we assume that $x$ follows a Gaussian distribution $\mathcal{N}_{0}=\mathcal{N}(\hat{\mu}_{k},\hat{\Sigma}_{k})$. According Matrix Cookbook [38] and [58], the following identity holds: $\mathbb{E}_{\mathcal{N}_{0}}\left[(x-\mu_{k})^{\top}\Sigma^{-1}(x-\mu_{k})\right]\!=\!(\hat{\mu}_{k}-\mu_{k})^{\top}\Sigma^{-1}(\hat{\mu}_{k}-\mu_{k})+\text{Tr}(\Sigma^{-1}\hat{\Sigma}_{k})$. Based on this, the $\mathrm{KL}$ divergence between the distributions $\mathcal{N}_{0}=\mathcal{N}(\hat{\mu}_{k},\hat{\Sigma}_{k})$ and $\mathcal{N}_{1}=\mathcal{N}(\mu_{k},\Sigma))$ is given as:
@@ Discussion.
Appendix B More Experimental Results
B.1 Experiments with CLIP-RN50 Backbone
@@
Finally, we evaluate the ADAPT ’s adaptation on fine-grained categorization benchmarks, as shown in Table 14. In the Online setting, ADAPT achieves 62.82% average accuracy, which is competitive with the strongest methods like DMN, while maintaining the advantage of being backpropagation-free. In the transductive setting, ADAPT achieves the highest accuracy of 64.64%, outperforming prior state-of-the-art methods such as TransCLIP and StatA. It ranks at the top in most datasets, demonstrating strong generalization to subtle inter-class differences. These results highlight the effectiveness and robustness of our framework under fine-grained and distribution-shifted conditions.
@@ B.2 Few-shot Adaptation
B.3 Evaluation with Different VLMs

### v8

@@ 4.2 Main Results
Online CLIP Radford et al. (2021) 66.74 - 8m 0.79
TPT Manli et al. (2022) 68.95 2.21 9h 45m 4.29
DiffTPT Feng et al. (2023) 70.30 3.56 > 20h 4.60
TDA Karmanov et al. (2024) 69.51 2.77 50m 0.84
TPS Sui et al. (2025) 70.38 3.64 1h 19m 1.71
@@
Trans. GDA-CLIP  Wang et al. (2024) 64.13 -2.61 1.31m 10.03
TransCLIP Zanella et al. (2024a) 70.30 3.56 1.34m 16.17
StatA Zanella et al. (2025) 69.90 3.16 1.5m 20.74
@@
Vanilla Radford et al. (2021) 64.70 27.52 67.43
Ensemble Radford et al. (2021) 66.56 28.54 67.74
CLIP Template Radford et al. (2021) 66.51 28.44 67.62
GPT  Zhu et al. (2024b) 66.53 28.56 70.76
@@ 4.3 Ablation Studies and Further Analysis
We study how different initializations of the mean $\hat{\boldsymbol{\mu}}$ in Eq. (9) affect performance. We compare 6 types of text-derived prototypes: CLIP’s vanilla template, 7 generic templates, 80 ImageNet-specific templates, GPT-generated descriptions in Zhu et al. (2024b), and its variants. As shown in Table 9, diverse and descriptive initializations improve performance, though ADAPT remains robust across all settings. We use GPT-based descriptions in our main results.
@@ A.2 Online Test-time Distribution Estimation
For the third term, we assume that $x$ follows a Gaussian distribution $\mathcal{N}_{0}=\mathcal{N}(\hat{\mu}_{k},\hat{\Sigma}_{k})$. According to the Matrix Cookbook [39] and [59], the following identity holds: $\mathbb{E}_{\mathcal{N}_{0}}\left[(x-\mu_{k})^{\top}\Sigma^{-1}(x-\mu_{k})\right]\!=\!(\hat{\mu}_{k}-\mu_{k})^{\top}\Sigma^{-1}(\hat{\mu}_{k}-\mu_{k})+\text{Tr}(\Sigma^{-1}\hat{\Sigma}_{k})$. Based on this, the $\mathrm{KL}$ divergence between the distributions $\mathcal{N}_{0}=\mathcal{N}(\hat{\mu}_{k},\hat{\Sigma}_{k})$ and $\mathcal{N}_{1}=\mathcal{N}(\mu_{k},\Sigma))$ is given as:
@@ Discussion.
Table 12: Comparison between ADAPT and related online TTA paradigms. We distinguish between transductive TTA, batch-based online TTA (Online, bs>>1), and the stricter single-instance online TTA (Online, bs=1). ADAPT supports all three protocols while remaining training/tuning-free.
Key Distinctions ADAPT (Ours) T3A [19] DPCore [67] DynaPrompt [55]
Core Paradigm Distribution-Aware Prototype-Based Prompt-Tuning Prompt-Tuning
Knowledge Bank Elements Confident Test Samples Learnable Prompts, Statistics Learnable Prompts
Training/Tuning Free $\times$ $\times$
Update Basis Sample-by-Sample Batch-Based Batch-Based Sample-Wise Tuning
Key Dependencies Batch of Data Batch of Data Augmentation, Hard Thresholds
Applicability (Trans.)
Applicability (Online, bs>>1) $\times$
Applicability (Online, bs=1) $\times$ $\times$
Comparison with Related Online TTA Paradigms.
We further compare ADAPT with representative test-time adaptation methods, including T3A [19], DPCore [67], and DynaPrompt [55]. Although these methods are often broadly referred to as online TTA, the term “online” may correspond to two different protocols. The first is single-instance online TTA (Online, bs=1), where test samples arrive one by one and must be processed without access to future or batched samples. The second is batch-based online TTA (Online, bs>>1), where each adaptation step can use a mini-batch of target samples and thus exploit batch-level statistics.
As summarized in Table 12, existing methods differ in their adaptation mechanisms and protocol assumptions. T3A and DPCore mainly rely on batch-level target information, making them more suitable for batch-based online or transductive settings. DynaPrompt can operate in the single-instance setting, but requires sample-wise prompt tuning and additional dependencies such as augmentation and hard thresholds. In contrast, ADAPT updates class-conditional statistics sample by sample using confident test samples, without gradient-based tuning. Therefore, ADAPT is applicable to both online protocols as well as the transductive setting, while remaining training/tuning-free and distribution-aware.
Appendix B More Experimental Results
Table 13: Stability analysis with CLIP-ViT-B/16 backbone. We report mean accuracy and standard deviation over 10 random permutations of the online test stream.
Task 1: Natural Distribution Shift
Method ImageNet ImageNet-A ImageNet-V ImageNet-R ImageNet-S OOD Avg. Avg.
ADAPT $70.72{\scriptstyle\pm 0.10}$ $63.45{\scriptstyle\pm 0.21}$ $64.56{\scriptstyle\pm 0.16}$ $80.55{\scriptstyle\pm 0.09}$ $53.08{\scriptstyle\pm 0.09}$ $5.41{\scriptstyle\pm 0.06} 66$$6.47{\scriptstyle\pm 0.06}$
Task 2: Corruption Robustness
Method Blur Weather
Defo. Glas. Moti. Zoom Snow Fros. Fog Brig.
ADAPT $26.34{\scriptstyle\pm 0.06}$ $17.90{\scriptstyle\pm 0.07}$ $27.26{\scriptstyle\pm 0.09}$ $25.42{\scriptstyle\pm 0.08}$ $36.25{\scriptstyle\pm 0.07}$ $34.66{\scriptstyle\pm 0.07}$ $40.89{\scriptstyle\pm 0.06}$ $60.28{\scriptstyle\pm 0.06}$
Digital Noise Avg.
Cont. Elas. Pix. JPEG Gauss. Shot Impu.
$19.84{\scriptstyle\pm 0.22}$ $16.10{\scriptstyle\pm 0.06}$ $37.30{\scriptstyle\pm 0.10}$ $37.23{\scriptstyle\pm 0.08}$ $15.71{\scriptstyle\pm 0.05}$ $16.82{\scriptstyle\pm 0.05}$ $15.81{\scriptstyle\pm 0.07}$ $28.52{\scriptstyle\pm 0.03}$
Task 3: Fine-Grained Categorization
Method Aircraft Caltech Cars DTD EuroSAT Flower Food101 Pets Sun397 UCF101 Avg.
ADAPT $28.93{\scriptstyle\pm 0.21}$ $94.60{\scriptstyle\pm 0.15}$ $68.05{\scriptstyle\pm 0.27}$ $53.87{\scriptstyle\pm 0.61}$ $67.24{\scriptstyle\pm 0.71}$ $76.28{\scriptstyle\pm 0.34}$ $83.91{\scriptstyle\pm 0.04}$ $90.37{\scriptstyle\pm 0.58}$ $71.26{\scriptstyle\pm 0.26}$ $71.82{\scriptstyle\pm 0.45}$ $70.64{\scriptstyle\pm 0.10}$
B.1 Stability Analysis of the Results.
To further evaluate the stability of ADAPT, we evaluate ADAPT over 10 random permutations of the online test stream, reporting the average performance and standard deviation across three tasks. As shown in Table 13, these results confirm our method’s high stability, with a consistently low standard deviation across all tasks.
@@ Appendix B More Experimental Results
B.2 Experiments with CLIP-RN50 Backbone
@@
Finally, we evaluate the ADAPT on fine-grained categorization benchmarks, as shown in Table 16. In the Online setting, ADAPT achieves 62.82% average accuracy, which is competitive with the strongest methods like DMN, while maintaining the advantage of being backpropagation-free. In the transductive setting, ADAPT achieves the highest accuracy of 64.64%, outperforming prior state-of-the-art methods such as TransCLIP and StatA. It ranks at the top in most datasets, demonstrating strong generalization to subtle inter-class differences. These results highlight the effectiveness and robustness of our framework under fine-grained and distribution-shifted conditions.
@@ B.3 Few-shot Adaptation
B.4 Evaluation with Different VLMs
@@ B.5 Additional Analysis
Table 18: Comparison between TDA [22] and our ADAPT under comparable prompt conditions. ADAPT consistently outperforms TDA across all prompt settings, indicating that its improvement mainly comes from distribution-aware adaptation rather than advanced prompt engineering.
Task Method Vanilla Ensemble CLIP Template GPT GPT & Ensemble GPT & CLIP Template
Task 1 TDA $63.92$ $66.01$ $65.93$ $65.92$ $65.94$ $65.96$
ADAPT (Online) $4.70{\scriptstyle~(+0.78)} 66$$56{\scriptstyle~(+0.55)} 66.$$51{\scriptstyle~(+0.58)} 66.$$53{\scriptstyle~(+0.61)} 66.$$57{\scriptstyle~(+0.63)} 66.$$6.58{\scriptstyle~(+0.62)}$
ADAPT (Trans.) $65.53{\scriptstyle~(+1.61)}$ $67.16{\scriptstyle~(+1.15)}$ $67.23{\scriptstyle~(+1.30)}$ $67.09{\scriptstyle~(+1.17)}$ $67.10{\scriptstyle~(+1.16)}$ $67.12{\scriptstyle~(+1.16)}$
Task 2 TDA $27.17$ $28.25$ $28.05$ $28.38$ $28.76$ $28.69$
ADAPT (Online) $27.52{\scriptstyle~(+0.35)}$ $28.54{\scriptstyle~(+0.29)}$ $28.44{\scriptstyle~(+0.39)}$ $28.56{\scriptstyle~(+0.18)}$ $28.98{\scriptstyle~(+0.22)}$ $28.91{\scriptstyle~(+0.22)}$
ADAPT (Trans.) $29.45{\scriptstyle~(+2.28)}$ $30.00{\scriptstyle~(+1.75)}$ $29.96{\scriptstyle~(+1.91)}$ $30.29{\scriptstyle~(+1.91)}$ $30.51{\scriptstyle~(+1.75)}$ $30.48{\scriptstyle~(+1.79)}$
Task 3 TDA $66.92$ $67.39$ $64.46$ $69.62$ $69.04$ $69.18$
ADAPT (Online) $67.43{\scriptstyle~(+0.51)}$ $67.74{\scriptstyle~(+0.35)}$ $67.62{\scriptstyle~(+3.16)}$ $70.76{\scriptstyle~(+1.14)}$ $69.95{\scriptstyle~(+0.91)}$ $69.90{\scriptstyle~(+0.72)}$
ADAPT (Trans.) $69.55{\scriptstyle~(+2.63)}$ $70.00{\scriptstyle~(+2.61)}$ $70.38{\scriptstyle~(+5.92)}$ $72.43{\scriptstyle~(+2.81)}$ $72.05{\scriptstyle~(+3.01)}$ $71.91{\scriptstyle~(+2.73)}$
Table 19: Comparison with AWT [72] and AWT without visual augmentations. AWT shows a clear performance drop when augmentations are removed, while ADAPT achieves strong performance without requiring visual augmentations.
Method ImageNet Aircraft Caltech101 Cars DTD EuroSAT Flower102 Food101 Pets Sun397 UCF101 Avg.
AWT 71.32 29.22 95.54 69.93 55.56 58.61 75.07 85.54 92.53 70.58 72.51 70.58
AWT 69.03 27.48 94.77 65.94 53.96 60.41 75.11 84.76 91.88 68.18 69.81 69.21
ADAPT (Online) 70.91 28.95 94.48 68.19 55.20 68.19 75.56 83.81 92.01 70.57 70.66 70.78
ADAPT (Trans.) 71.56 30.81 95.46 71.32 56.86 65.93 80.11 85.15 92.59 72.25 73.86 72.35
Further Comparison with TDA under Comparable Prompt Conditions.
We further compare ADAPT with TDA [22] under comparable prompt conditions to clarify that ADAPT’s performance gains mainly stem from its distribution-aware adaptation mechanism rather than advanced prompt engineering. TDA, as a representative memory-based method, relies on a non-parametric, instance-wise KNN-like strategy, which can make predictions sensitive to individual noisy samples in online scenarios. In contrast, ADAPT adopts a parametric Gaussian discriminant analysis (GDA)-based strategy to explicitly model class-conditional feature distributions. This formulation reduces the influence of individual outliers and leads to more stable decision boundaries.
For a fair comparison, we evaluate ADAPT and TDA using the same set of prompting strategies, including vanilla prompts, ensemble prompts, CLIP templates, GPT prompts, GPT ensemble prompts, and GPT & CLIP templates. As shown in Table 18, ADAPT consistently outperforms TDA across all tasks and prompt settings. These results indicate that the improvements of ADAPT primarily come from its distribution-aware adaptation mechanism rather than stronger textual prompts. Moreover, the prior-removal analysis in the main paper shows that ADAPT maintains strong performance even when the prompt-based CLIP prior (Vanilla) is removed, further supporting that its robustness mainly comes from data-driven distribution modeling.
Further Comparison with AWT.
While our ADAPT uses the same handcrafted prompts as AWT [72], their core adaptation strategies differ fundamentally. AWT aligns multiple augmented views of each test image with textual prompts via Optimal Transport, making its performance highly dependent on visual augmentations. To illustrate this distinction, we compare ADAPT with both the original AWT and a variant without visual augmentations, denoted as AWT. As shown in Table 19, AWT suffers a significant performance drop without augmentations, confirming its reliance on augmented visual views. In contrast, ADAPT achieves consistently strong performance without requiring such augmentations. These results highlight ADAPT’s ability to perform source-free and training-free adaptation without relying on computationally expensive visual augmentations. This makes ADAPT a more streamlined and versatile solution that is readily applicable to both online and transductive settings.
@@
Table 22: Ablation study on explicit memory bank pruning under the online protocol.
Method ImageNet ImageNet-A
ADAPT (Online) 70.91 63.32
w/ 0.1 pruning 69.16 50.69
w/ 0.3 pruning 69.03 50.60
w/ 0.5 pruning 69.02 50.56
w/ 0.7 pruning 68.97 50.52
Effect of Explicit Memory Bank Pruning.
We further investigate whether explicitly pruning low-confidence samples from the memory bank can improve the robustness of ADAPT. Specifically, we periodically prune the memory bank every 1k test samples by removing a fixed portion of the lowest-confidence samples from each class. We consider pruning ratios of 0.1, 0.3, 0.5, and 0.7. The pruning is performed on a per-class basis, since applying a single global confidence threshold may unfairly remove samples from naturally harder classes.
Table 22 reports the results on ImageNet and ImageNet-A under the online protocol. We observe that explicit bank pruning consistently degrades performance, especially on ImageNet-A, which contains a stronger distribution shift. For example, pruning only 10% of the lowest-confidence samples reduces the ImageNet-A accuracy from 63.32% to 50.69%, and more aggressive pruning leads to similar or slightly worse performance. These results suggest that ADAPT already provides a form of soft dynamic filtering through its regularized probabilistic estimation. Rather than discarding samples entirely, ADAPT naturally reduces the influence of unreliable samples during distribution estimation. In contrast, explicit pruning may prematurely remove samples that are initially uncertain but could become informative as adaptation progresses. Therefore, we keep the original memory-bank design without explicit pruning, which provides better stability and robustness under distribution shift.
Table 23: Robustness under early noisy predictions in long-tailed adaptation.
Method Head class Med class Tail class Avg.
CLIP 68.44 69.69 68.42 69.33
CLIP w/ Noise $62.16{\scriptstyle~(-6.28)}$ $64.03{\scriptstyle~(-5.66)}$ $62.09{\scriptstyle~(-6.33)}$ $63.25{\scriptstyle~(-6.08)}$
ADAPT (Online) 69.22 70.31 69.08 69.83
ADAPT w/ Noise $63.55{\scriptsize~(-5.67)}$ $65.20{\scriptsize~(-5.11)}$ $62.25{\scriptsize~(-6.83)}$ $64.27{\scriptsize~(-5.56)}$
Long-tailed Adaptation with Early Noisy Predictions.
We further evaluate ADAPT under a challenging long-tailed scenario with noisy early-stage predictions. Specifically, we construct a long-tailed subset from ImageNet-1K by sampling classes according to a power-law distribution, resulting in 61 head classes ($\geq 40$ shots, 2,719 images), 338 medium classes (10–40 shots, 7,396 images), and 601 tail classes ($\leq 10$ shots, 2,416 images). To simulate early memory contamination, we inject noise during the first 30% phase of adaptation by randomly reassigning 30% of the high-confidence predictions to incorrect classes.
As shown in Table 23, while both CLIP and ADAPT degrade under this adverse setting, ADAPT consistently achieves higher final accuracy than CLIP (64.27% vs. 63.25%), demonstrating stronger overall robustness. For the extremely few-shot tail classes, both methods are highly vulnerable to the strong initial misguidance and exhibit a comparable performance drop. This is expected, as a 30% initial corruption can heavily skew parameter estimates when very few samples are available. Nevertheless, ADAPT shows a stronger self-correction capability as adaptation progresses. The accumulation of correctly identified samples, especially from head and medium classes, allows ADAPT to progressively refine its distribution estimates and mitigate the influence of initial noisy predictions. These results indicate that ADAPT is more robust than the CLIP baseline under noisy long-tailed adaptation, although extremely few-shot tail classes remain challenging.